VPTOVF(1) VPTOVF(1) NAME vptovf−convertvirtualpropertyliststovirtualfontmetrics SYNOPSIS vptovf[−verbose]vplname[.vpl]vfmname[.vf]tfmname[.tfm] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thevptovfprogramtranslatesa(human-oriented)propertylistfiletoapairof(program-ori-ented)filesintheVF(virtualfont)andTFM(TEXfontmetric)formats.Thus,avirtualfontfile canbeeditedanditsexactcontentscanbedisplayedmnemonically.Newvirtualfonts,which mapcharactersasseenbyTEXintoanarbitrarysequenceoflow-leveltypesettingoperations,can alsobecreatedinthisway. Allthreefilenames,vplname,vfname,andtfmname,areextendedwiththeappropriatesuffix, iftheylackone. OPTIONS Withoutthe−verboseoption,vptovfoperatessilently.Withit,abannerandprogressreport areprintedonstdout. SEEALSO pltotf(1),tftopl(1),vftovp(1). AUTHORS DonaldE.Knuthwrotetheprogram,basedinpartonanideaofDavidFuchs,startingwiththe codeforpltotf(1).KarlBerryadapteditforcompilationwithweb2c. Web2C7.32x 16December1994 1